OpenStudy (anonymous):

May someone please explain to me how to do this Algebra problem? Write an equation of the line that passes through each point with the given slope. (-5, 5), m=-2 I have no idea.... I know about y=mx+b So I presume it'll start off with y=-2x, but after that, I don't know. Any assistance will be greatly appreciated. Thank you very much.

OpenStudy (anonymous):

Funny. In school here in the UK, we use y= mx + c. Ok, m is the gradient and here it's -2 so it's negative and therefore downward sloping. It's absolute magnitude is 2 meaning that for every change of 1 in x there is a change of 2 in y - so it's a steep downward sloping line which intersects at the point where y = 5 , and that represents the c in the equation y = mx + c, or in your case, the b in y = mx + b

OpenStudy (pfenn1):

Just for future reference, this question would be better suited to the Mathematics subject, not Writing.